OpenStudy (anonymous):

solve by graphing 20x+40y=240.00 10x+80y=300.00

OpenStudy (anonymous):

@phi

OpenStudy (phi):

this is from your previous problem? Because the numbers are so big, I would simplify both equations first (divide the first by 20 and the second by 10)

OpenStudy (anonymous):

I got x+2y=12 x+8y=30

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Are you supposed to graph these lines on paper, or with a computer ?

OpenStudy (anonymous):

on paper

OpenStudy (phi):

remember, you already know the answer (3 and 6 ? I forget which is x and which is y) can you refresh my memory?

OpenStudy (anonymous):

x is senior

OpenStudy (anonymous):

x is 6

OpenStudy (anonymous):

y is 3 @phi

OpenStudy (phi):

that is where the lines will cross (this is cheating because we would not know that in advance). x+2y=12 when x is 0, what is y ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@phi

OpenStudy (phi):

0 + 2y= 12 or 2y= 12 I would divide both sides by 2

OpenStudy (anonymous):

i thought x would be 1

OpenStudy (anonymous):

if you divide both sides by 2 it will be 6

OpenStudy (phi):

yes. The idea is x+2y=12 is a line and that means for each x value, there is a y value i..e you get an (x,y) point for example when you make x=0 you get 0+2y =12 2y= 12 y= 6 (after dividing both sides by 2) so (0,6) is on the line can you plot that point ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

if you choose x=1 we get 1+2y= 12 2y = 11 (after add -1 to both sides) y= 11/2 that is a pain to plot. Try x= 2 2+2y= 12 can you find y ?

OpenStudy (anonymous):

5

OpenStudy (phi):

so (2,5) is another point plot that point. then draw a line through the two points and as a check , remember this line should also go through (6,3)

OpenStudy (anonymous):

why would 2,5 be another point

OpenStudy (anonymous):

why are we using 2,5 i thought we had to solve x+8y=30

OpenStudy (arcticfoxzx87):

can i also help

OpenStudy (anonymous):

sure

OpenStudy (phi):

we are still doing the first line. to plot a line, you need two points we found two points on the first line.

OpenStudy (arcticfoxzx87):

when i do it it ends up as a slope intercept dont know if its right

OpenStudy (anonymous):

ooo ok @phi

OpenStudy (phi):

once you plotted the two points, draw a line through them (and be sure the line goes thought (6,3)

OpenStudy (arcticfoxzx87):

brb

OpenStudy (anonymous):

okay it does

OpenStudy (phi):

now we do the 2nd line

OpenStudy (phi):

x+8y=30

OpenStudy (anonymous):

80y=300.00

OpenStudy (anonymous):

(0,3.75)

OpenStudy (phi):

yes, but it's hard to plot fractions also, it's easier to use the equation after we simplified it (notice if we use x+8y=30 with x=0 we get 8y= 30 y= 30/8 = 3.75

OpenStudy (phi):

x+8y=30 if we make y=0 we get x=30 so (30,0) is on the line but that point is a bit far away ... it depends how big your graph paper is but if we let y= 2, we get x+8*2= 30 and x+16=30 x= 14 so (14,2) is a good point try y = 4 in x+8y= 30

OpenStudy (anonymous):

x+32=30

OpenStudy (phi):

ok , now add -32 to both sides

OpenStudy (anonymous):

2

OpenStudy (anonymous):

-2

OpenStudy (phi):

yes, x=-2 so (-2,4) is a point. can you plot (-2,4) and (14,2) and draw a line through the points ? (and be sure it goes through (6,3)

OpenStudy (phi):

The final result should look like this

OpenStudy (anonymous):

thank you @phi